OpenStudy (anonymous):

Solve the system using the substitution method. Show your work. Check your solution. 4y + 2x = 48 x = –2y + 26

OpenStudy (anonymous):

in substitution method we substitue value of any variable in terms of another here.. 4y + 2x = 48 2x = 48 - 4y x = (48 - 4y)/2 we substitue this value of x in second eqn

OpenStudy (anonymous):

wait but we havents found x

OpenStudy (anonymous):

x=11

OpenStudy (anonymous):

show ur work..n pls check the qstn once...i guess sumthng is wrong

OpenStudy (anonymous):

no solution!

OpenStudy (anonymous):

i did it wrong?

OpenStudy (anonymous):

yes ... i am also getting no solution

OpenStudy (anonymous):

my tutor said it might not have a solution so we are right!

OpenStudy (anonymous):

for this pair of eqn there is no solution ...iam 100 % sure

OpenStudy (anonymous):

we are right then! THANK YOU !

OpenStudy (anonymous):

welcome..any other questions

OpenStudy (anonymous):

yes Line A goes through the points (4, 0) and (–2, 18). Line B goes through the points (0, –5) and (2, –11). Are the two lines parallel, perpendicular, or neither?

OpenStudy (anonymous):

i got neither

OpenStudy (anonymous):

find the slope of both the lines for line A slope is -3 and for line B slope is also -3 so both the lines are parallel

OpenStudy (anonymous):

because they have the same slope ?

OpenStudy (anonymous):

yes m1 = m2 then parallel and when m1m2 + -1 en perpndclr

OpenStudy (anonymous):

its perpendicular when they are opposites?

OpenStudy (anonymous):

no perpendicular when m1m2 = -1 or m1 = -1/m2

OpenStudy (anonymous):

if its a horizontal line the slope is 0 right?

OpenStudy (anonymous):

because there is no rise over run

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i have one more question if you would like to go through it with me

OpenStudy (anonymous):

im reviewing for a test

OpenStudy (anonymous):

yes sure...

OpenStudy (anonymous):

Solve the system of equations by addition method. Show your work. Check your solution. 3𝑎+2𝑏=11 7𝑎−3𝑏= −5

OpenStudy (anonymous):

i dont know if those are zeros or boxes